A staircase has steps each of height 20cm and width 40 cm. A ball rolls off the top of the staircase horizontally with velocity 5m/s. The ball hits the nth step from the top at the first bounce, n=?a) 5 b)6 c)7 d)8

Grade:12

1 Answers

Rama sai rohan
6 years ago
Range =n×0.4Height =n×0.2Formula:R=u√(2h/g) n×0.4=5×√(2×0.2n/g) If we solve this we get n=6.2 so, we take it as 6th step